½(y-2)+2 =3/8(3y-4)

I am so confused with this problem. can you show me how to solve it step by step.
Thank you
Michelle

Answer by ewatrrr(10682)


Hi,
***Note: let's proceed to have all denominators = 1
½(y-2)+ 2 =3/8(3y-4)
multiplying each term on both sides of the equation by 8 so as all denominators = 1
4(y-2) + 2*8 = 3(3y-4)
4y - 8 + 16 = 9y - 12
-8 + 16 + 12 = 5y
20 = 5y
4 = y
checking our answer
½(4-2)+ 2 =3/8(3*4-4) = 3